Apply second substitution theorem evaluate


i) integral 1 to 9 (√ t)/(2+√t)





1
Expert's answer
2022-03-27T07:26:30-0400
"\int \frac{\sqrt{t}}{\sqrt{t}+2} \mathrm{~d} t\\[2mm]
\text{Substitute } u=\sqrt{t}+2 \longrightarrow \frac{\mathrm{d} u}{\mathrm{~d} t}=\frac{1}{2 \sqrt{t}}\longrightarrow \mathrm{d} t=2 \sqrt{t} \mathrm{~d} u, \text{ use:}\\
\begin{aligned}
& t=(u-2)^{2} \\
=& 2 \int \frac{(u-2)^{2}}{u} \mathrm{~d} u
\end{aligned}
\\[4mm]
\text{Now solving:}\\[2mm]
\begin{aligned} & \int \frac{(u-2)^{2}}{u} \mathrm{~d} u \\=& \int\left(u+\frac{4}{u}-4\right) \mathrm{d} u \end{aligned}\\[2mm]
\text{Apply linearity:}\\[2mm]
=\int u \mathrm{~d} u+4 \int \frac{1}{u} \mathrm{~d} u-4 \int 1 \mathrm{~d} u\\[3mm]
\text{Now solving:}\\[2mm]
\int u \mathrm{~d} u\\[2mm]
\text{Apply power rule:}\\[2mm]
\begin{aligned}
\int u^{\mathrm{n}} \mathrm{d} u=& \frac{u^{\mathrm{n}+1}}{\mathrm{n}+1} \text { with } \mathrm{n}=1 \\
&=\frac{u^{2}}{2}
\end{aligned}
\\[2mm]"

"\text{Now solving:}\\[2mm]
\int \frac{1}{u} \mathrm{~d} u\\[2mm]
\text{This is a standard integral:}\\[2mm]
=\ln (u)\\[3mm]
\text{Now solving:}\\[2mm]
\int 1 \mathrm{~d} u\\[2mm]
\text{Apply constant rule:}\\[2mm]
=\boldsymbol{u}\\[2mm]
\text{Plug in solved integrals:}\\[2mm]
\begin{gathered}
\int u \mathrm{~d} u+4 \int \frac{1}{u} \mathrm{~d} u-4 \int 1 \mathrm{~d} u \\
=4 \ln (u)+\frac{u^{2}}{2}-4 u
\end{gathered}\\[2mm]
\text{Plug in solved integrals:}\\[2mm]
2 \int \frac{(u-2)^{2}}{u} \mathrm{~d} u\\=8 \ln (u)+u^{2}-8 u\\[2mm] \text{ Undo substitution }\\ u=\sqrt{t}+2 =-8(\sqrt{t}+2)+(\sqrt{t}+2)^{2}+8 \ln (\sqrt{t}+2)\\[2mm]
\text{The problem is solved:}\\[2mm]
\begin{gathered}
\int \frac{\sqrt{t}}{\sqrt{t}+2} \mathrm{~d} t \\
=-8(\sqrt{t}+2)+(\sqrt{t}+2)^{2}+8 \ln (\sqrt{t}+2)+C
\end{gathered}""=t-4 \sqrt{t}+8 \ln (\sqrt{t}+2)+C"

Evaluating the above function at [1,9], we have:


"8 \ln (5)-8 \ln (3)\\[3mm]

\text{Simplify/rewrite:}\\[2mm]
8(\ln (5)-\ln (3))\\[3mm]
\text{ Approximation:}\\4.086604990127925"